Rhodo132:

(there is no graph) Relationship 1 has an initial value of 5 and a rate of change of 3/2. Relationship 2 has a rate of change of 5/6. Explain how to find what the initial value of Relationship 2 needs to be for it to have an output of 50 with the exact same input as Relationship 1.

Angle:

Hello and Welcome :) do you think you can put these equations into formats such as y = mx + b? for example, m = rate of change b = initial value y = output x = input so for Relationship 2, output = y = 50 rate of change = m = 5/6 so we would write 50 = (5/6)x + b does this make sense?

Rhodo132:

ha, not at all. I suck at math... I can try though